A particle starts moving at t=0 from origin along positive X axis. Its speed depends on the distance traveled x as v=k√x . If v=dx/dt and a=dv/dt, then find x,v and a in terms of time.
The answer is k(square) t(squ)/4, k(squ)t/2, k(squ)/2 respectively.
pls explain

Grade:11

1 Answers

Khimraj
5 years ago
v = k√x
dx/dt = k√x
dx/√x = kdt
integerating both side
2√x = kt
x = k2t2/4
v = dx/dt = k2t/2
a = dv/dt = k2/2
Hope it clears.............................................